Use substitution to solve: 4x-27=-28 and 2x+3y=-6
Answer by Emsley(35)

From 2 2x=-6-3y sub in 1 -12-6y-27=-28 -6y=-28+27+12-1=10 y=10/-6=-5/3
sub in 1 to get x